Assume 30-year maturity that interest is 4.00% for the market rate of interest and a mortgage balance of $110,000.00. What will the balance of the mortgage be after 6.00 years

a. $1,194,584.03

b. $525.16

c. $97,127.02

d. $364,484.78
Answer #1

Balance of the mortgage after 6 years=Loan*(1+rate/12)^(12*6)-Loan/(1-1/(1+rate/12)^(12*loan term))*((1+rate/12)^(12*6)-1)=110000*(1+4%/12)^(12*6)-110000/(1-1/(1+4%/12)^(12*30))*((1+4%/12)^(12*6)-1)=97127.02301